OpenStudy (anonymous):

Enter the correct solution to the system of equations below as an order pair.(x,y) y=x+2 y=-2x+2

OpenStudy (callisto):

y=x+2 -(1) y=-2x+2 -(2) (1)-(2) y-y=x+2 - (-2x+2 ) 0 = 3x x=0. Put x =0 into (1), what would you get?